1C2H4+3O2—> 2CO2+2H2O is you start with 45 g of (C2H4), how many g of Carbon Dioxide will be produced?
1
Expert's answer
2020-01-24T05:00:49-0500

The balanced equation is :

"C_2H_4+3O_2\to 2CO_2+2H_2O"

From the above equation we can see that the ratio of stochiometric coefficient of "C_2H_4 \ to \ CO_2" is 1:2

which means from 1 moles of "C_2H_4" 2 moles of carbon dioxide will be produced

we are given 45 grams of "C_2H_4" converting them into number of moles

Number of moles = given mass/ molar mass = 45/48=0.9375

moles of carbon dioxide produced will be equal to twice of this

moles of carbon dioxide = 2*0.9375 = 1.875

Converting this into mass of carbon dioxide produced = 1.875*44 = 82.5 grams